# Tail calibration of probabilistic forecasts

### v2

@@ 1 Introduction
In some fields, forecast calibration refers to re-calibration methods that adapt forecasts so that they are calibrated. This paper does not concern re-calibration, though it does facilitate the introduction of tail re-calibration methods. Tail re-calibration methods currently do not exist, largely due to the lack of tools available to evaluate the resulting forecasts. Our definitions of tail calibration remedy this by allowing forecasters to assess whether their forecasts for extreme outcomes are reliable, and if not, to understand how the forecasts are miscalibrated, for example, if the predictive distribution has tails that are too light or too heavy.
In Section 2, we revisit notions of calibration for probabilistic forecasts, and introduce concepts that are suitable when evaluating forecasts for extreme events. Connections between tail calibration and well-known results in extreme value theory are explored in Section 3. Section 4 presents several examples and introduces diagnostic tools to assess tail calibration in practice, and these diagnostic tools are then applied to a canonical weather forecasting data set in Section 5. Section 6 summarizes our findings. The appendices in the supplement contain sections about the following topics: additional proofs and calculations; a result about the evaluation of tail behavior using multiple scoring rules; a definition of, and results on, marginal tail calibration; and additional simulation results evaluating well-known forecasts in the literature with respect to their probabilistic tail calibration. R code to reproduce the results herein is available at https://github.com/sallen12/TailCalibration.
@@ 2 Tail calibration
Let $(\Omega,\mathcal{A},\operatorname{\mathsf{P}})$ be a probability space on which a random pair $(F,Y)$ is defined, where $F$ is a random probabilistic forecast for the outcome variable $Y$. We restrict attention to real-valued outcomes and treat $F$ as a random cumulative distribution function (cdf). We assume that $F$ is almost surely continuous, as this case is most relevant with regards to extremes and leads to a simpler presentation. However, generalizations are possible; see Remark 1.
@@ 2.2 Notions of tail calibration
To assess the calibration of forecasts for extreme events, we adopt the standard approach of extreme value theory and decompose $\operatorname{\mathsf{P}}(Y>x+t)=\operatorname{\mathsf{P}}(Y>t)\operatorname{\mathsf{P}}(Y>x+t\mid Y>t)$, for $x\geq 0$ and large $t$, i.e., $\operatorname{\mathsf{P}}(Y>t)$ positive but close to zero. Then, we separately consider calibration with respect to both factors on the right-hand side. Under auto-calibration, we have
@@
To avoid this scaling issue , we define the excess probability integral transform (excess PIT) as
@@ 3 Probabilistic tail calibration and peaks-over-threshold models
    The distribution of $Y$ is in the max-domain of attraction of a non-degenerate distribution.
Assumption (A1) is standard in extreme value theory, and by Balkema and de Haan (1974), it is equivalent to the assumption that there exist $\xi\in\mathbb{R}$ and a scaling function $\sigma_{Y}(\,\cdot\,)>0$ such that
@@
For the proof of Proposition 14, see the more general Proposition 31 in Appendix C in the supplement. If the scale parameter $\sigma_{F}(t)$ in (9) is allowed to be random, then $F$ can be probabilistically tail calibrated for $Y$ even when the generalized Pareto shape parameters $\xi$ and $\eta$ in (8) and (9) differ. For instance, if $\Delta\sim\Gamma(1/\gamma,1/\gamma)$ for some $\gamma>0$ and if, conditionally on $\Delta$, the distribution of $Y$ is $F=\operatorname{Exp}(\Delta)$, i.e., $\operatorname{\mathsf{P}}(Y\leq x\mid\Delta)=1-e^{-\Delta x}=F(x)$ for $x\geq 0$, then the unconditional distribution of $Y$ is $\operatorname{GPD}_{1,\gamma}$, so that (8) holds with $\xi=\gamma$, while $F_{t}$ is $\operatorname{Exp}(\Delta)$ too, and thus $F_{t}(x/\Delta)=1-e^{-x}$, which is (9) with random scale parameter $\sigma_{F}(t)=1/\Delta$ and with $\eta=0$.
Assumptions (A1), (A2), and property (a) of Proposition 14 are sufficient but not necessary conditions for probabilistic tail calibration. Indeed, according to Proposition 7, for non-random probabilistic forecasts, it is sufficient that $\operatorname{\mathsf{P}}(Y>t)/(1-F(t))\to 1$ as $t\to x_{Y}$.
@@ 4.1 Probabilistic tail calibration
Figure 5 displays the diagnostic plots for the tail calibration of the three forecasters using $n=10^{6}$ independent realizations of $\Delta$ and $Y$, see also Figure 10. The extremist forecast clearly has a heavier tail than the true distribution of the outcomes, and the nature of this miscalibration does not change as the threshold is increased.
@@ 4.2 Tail $\mathcal{B}$-calibration
To assess tail $\mathcal{B}$-calibration, the forecast-observation pairs $(F_{1},y_{1}),\dots,(F_{n},y_{n})$ can be stratified into $J\in\mathbb{N}$ mutually disjoint bins, $B_{1},\dots,B_{J}$, depending on variables that represent the information contained in $\mathcal{B}$. The diagnostic plots for probabilistic tail calibration can then be applied separately to each bin of the forecasts and observations; if the forecasts appear probabilistically calibrated for all bins, then there is evidence to suggest that they are tail $\mathcal{B}$-calibrated. For example, the combined ratio in (10) can be calculated for each bin,
@@
Displaying the diagnostic plots for each forecast method and each bin leads to a large number of plots to analyze. To simplify visualization, (tail) mis-calibration can be summarized in a single value. This aligns with the general framework to assess conditional calibration proposed by Tsyplakov (2011) based on moment conditions and test functions, which is akin to how forecasts are compared in Giacomini and White (2006) and Nolde and Ziegel (2017). As a measure of tail calibration, consider the supremum distance $\sup_{u\in[0,1]}|\hat{R}_{t}(u)-u|$, which quantifies the maximum absolute difference between the combined ratio and the diagonal line in the diagnostic plot proposed above. Under probabilistic tail calibration, this difference should be close to zero for large $t$. To assess tail $\mathcal{B}$-calibration, $\hat{R}_{t}$ can be replaced by $\hat{R}_{t,j}$, and this distance can be calculated for $j=1,\dots,J$.
@@
for the ideal, climatological and extremist forecasters in Example 17 are shown as a function of the threshold in Figure 6, for $J=3$ bins for $\Delta$. For the ideal forecaster, the difference between $\hat{R}_{t,j}(u)$ and $u$ is essentially zero for all bins, confirming that it is tail $\sigma(\Delta)$-calibrated. The climatological forecaster, on the other hand, is clearly not calibrated conditionally on $\Delta$, despite being probabilistically tail and tail auto-calibrated. These plots for conditional calibration are shown for the empirical analogue (12) of the combined ratio in (4), though an analogous approach can be employed to evaluate the occurrence ratio (5) or the severity ratio (6).
@@ 5 Application
Three forecasting methods are compared. Firstly, we consider the 51-member ensemble forecasts generated by the ECMWF’s IFS ensemble. Ensemble forecasts are collections of point forecasts, forming an empirical predictive distribution function. An empirical distribution comprised of 51 sample values is unlikely to be useful when predicting extremes; in particular, the resulting forecast will assert that high thresholds will be exceeded with probability zero. We therefore also assess the performance of a smoothed version of the discrete ensemble forecast, which converts the IFS ensemble to a continuous forecast distribution by assuming that precipitation observations follow a censored logistic distribution, with location and scale parameters that are equal, respectively, to the mean and standard deviation of the corresponding IFS ensemble members. Finally, we present results for a statistical post-processing model that aims to remove systematic biases in the IFS ensemble to yield calibrated forecasts. The post-processing method also assumes that precipitation follows a censored logistic distribution, but with location and scale parameters that are affine functions of the IFS ensemble mean and standard deviation, respectively. This simple approach is generally referred to as ensemble model output statistics in the post-processing literature (Gneiting et al., 2005), and is commonly employed in studies of precipitation forecasts (see e.g. Messner et al., 2014). The parameters of the post-processing model are estimated by minimizing the continuous ranked probability score (CRPS) over 20 years of twice-weekly reforecasts at the corresponding station and lead time; further details regarding the data can be found in Demaeyer et al. (2023).
Figure 8 displays the probabilistic (tail) calibration diagnostic plots for the three forecast methods. For the discrete ensemble forecasts, the diagnostics for tail calibration are adapted as described in Remark 1. Results are shown at thresholds of 5, 10, and 15mm, roughly corresponding to the 97th, 99th, and 99.5th percentiles of the previously observed precipitation measurements (across all stations). Analogous conclusions are drawn when separate thresholds are employed at each station, corresponding to high quantiles of the local climatology, see Appendix E in the supplement. The raw IFS ensemble forecasts are under-dispersed and positively biased, whereas post-processing yields considerably more reliable forecasts. However, all forecasts are miscalibrated when focus is placed on extreme precipitation; the diagnostic plots indicate that all forecasts have a tail that is too light. This becomes more severe as the threshold of interest increases, suggesting that the forecasts are not probabilistically tail calibrated. Smoothing the IFS ensemble unsurprisingly improves tail calibration, but appears to worsen the overall calibration of the forecasts in this example. While post-processing improves upon the calibration of the raw ensembles, the resulting forecasts are still acutely miscalibrated when predicting extreme outcomes.
It is not trivial to derive a unified test for tail calibration based on the combined ratio at (10), and this becomes yet more difficult considering the limit as $t\to x_{Y}$. When assessing calibration with respect to a finite threshold, Mitchell and Weale (2023) propose separately testing whether the predicted threshold exceedance is correct, and whether the excess PIT values resemble a sample from a standard uniform distribution. Here, we implement a binomial test for the occurrence ratio, and a Kolmogorov–Smirnov test for uniformity of the excess PIT values, assuming they are independent and identically distributed. For thresholds 10mm and 15mm, all methods return a p-value of zero (up to numerical precision) on both tests, providing evidence against the forecasts being tail calibrated.
@@
The poor tail calibration of the post-processed forecasts is a deficiency of the method used, likely due to the choice of parametric distribution. This is irrespective of the number of ensemble members: Even if we had no covariate information, we could still issue the climatological distribution of the outcome, which is auto-calibrated and therefore tail auto-calibrated. However, when employing existing forecast evaluation techniques, practitioners would generally conclude that the post-processing method in this example generates reliable forecasts, which is misleading. In contrast, our diagnostic tools for tail calibration inform us that the forecast tails are too light, resulting in unreliable forecasts for extreme events. This information can be used to improve forecasts, for example by experimenting with alternative post-processing methods, based on different distributional assumptions.
Figure 8: Probabilistic tail calibration plots for the IFS and statistically post-processed forecast distributions in Section 5. Left: combined ratio (10); middle: pp-plot of $z_{i}^{t}$ for $i\in\mathcal{I}_{t}$; right: occurrence ratio (11) as a function of $t$. Results are shown for three thresholds, 5, 10, and 15mm, roughly corresponding to the 97th, 99th, and 99.5th percentiles of the previously observed precipitation measurements. The symbol ‘S’ denotes the standard notion of probabilistic calibration (i.e. for $t=-\infty$).
@@ 6 Discussion
Tail calibration can be assessed with respect to different information sets, though we imagine the unconditional case will be of most interest in practical applications. To assess probabilistic tail calibration, we introduce graphical diagnostic tools that not only demonstrate whether or not a forecast is tail calibrated, but also elucidate what types of biases occur in the tail; for example, whether the predictive distributions are too heavy- or light-tailed. Formal testing procedures for tail calibration are an interesting avenue for future work, where the crux will be to choose a data-dependent threshold that is representative of the asymptotic regime. Such tests could either be classical tests for fixed pre-specified sample sizes, or monitoring procedures as introduced by Arnold and Ziegel (2023), who propose a sequential procedure for testing probabilistic calibration based on $e$-processes, recognizing that forecasting is often an inherently sequential task.
@@
When the unconditional distribution of the outcome variable is in the max-domain of attraction of a non-degenerate distribution, i.e., satisfying Assumption (A1), probabilistic tail calibration is equivalent to correctly forecasting the shape and scale parameters of the GPD. Checks for probabilistic tail calibration therefore serve as a more general framework with which to validate tail models, since they are also applicable to distributions not satisfying Assumption (A1). Future work may investigate whether notions of tail calibration can conversely be used to estimate quantities relevant for extreme value analysis, such as tail indices.
@@ Appendix A Omitted proofs and calculations
Write $t_{n}=x_{Y}-1/n$. We have $\mathbf{1}\{Y>t_{n}\}\to\mathbf{1}\{Y\geq x_{Y}\}=\mathbf{1}\{Y=x_{Y}\}$. By the conditional dominated convergence theorem (e.g. Kallenberg, 2002, Exercise 8 on page 117), $\operatorname{\mathsf{P}}(Y>t_{n}\mid\mathcal{B})\to\operatorname{\mathsf{P}}(Y=x_{Y}\mid\mathcal{B})$ as $n\to\infty$ a.s. Since $\operatorname{\mathsf{E}}[\operatorname{\mathsf{P}}(Y=x_{Y}\mid\mathcal{B})]=\operatorname{\mathsf{P}}(Y=x_{Y})>0$, we have $\operatorname{\mathsf{P}}(B)>0$ where $B=\{\operatorname{\mathsf{P}}(Y=x_{Y}\mid\mathcal{B})>0\}$. Since $B\in\mathcal{B}$, we have
@@
$\frac{d}{du}\operatorname{\mathsf{P}}(Z\leq u)$ $=\frac{d}{du}\frac{1}{2}\{\operatorname{\mathsf{P}}(G_{+}(Y)\leq u)+\operatorname{\mathsf{P}}(G_{-}(Y)\leq u)\}$
@@
it follows that $z_{-}-1=z_{+}$, since $z\mapsto G(z)+G(z+1)$, $z\in[a-1,b]$ is strictly increasing, and $z_{+},z_{-}-1\in[a-1,b]$. Inserting this equality above yields $(d/du)\operatorname{\mathsf{P}}(Z\leq u)=1$, as required.

### v3

# supplement2

@@ 1 Introduction
In some fields, forecast calibration refers to re-calibration methods that adapt forecasts so that they are calibrated. This paper does not concern re-calibration, though it facilitates the introduction of tail re-calibration methods. Tail re-calibration methods currently do not exist, largely due to a lack of tools to evaluate the resulting forecasts. Our definitions of tail calibration remedy this by allowing forecasters to assess whether their forecasts for extreme outcomes are reliable, and if not, to understand how the forecasts are miscalibrated, for example, if the tails of the predictive distribution are too light or too heavy.
In Section 2, we revisit notions of calibration for probabilistic forecasts, and introduce concepts that are suitable when evaluating forecasts for extreme events. Connections between tail calibration and well-known results in extreme value theory are explored in Section 3. Section 4 presents several examples and introduces diagnostic tools to assess tail calibration in practice, and these diagnostic tools are then applied to a canonical weather forecasting data set in Section 5. Section 6 summarizes our findings. The appendices in the supplement contain sections about the following topics: additional proofs and calculations; a result about the evaluation of tail behavior using multiple scoring rules; a definition of, and results on, marginal tail calibration; additional simulation results evaluating well-known forecasts in the literature with respect to their probabilistic tail calibration; and explanations on the construction of confidence intervals in our diagnostic tools. An R package and R code to reproduce the results herein is provided as supplementary material.
@@ 2 Tail calibration
Let $(\Omega,\mathcal{A},\operatorname{\mathsf{P}})$ be a probability space supporting a random pair $(F,Y)$, where $F$ is a random probabilistic forecast for the outcome variable $Y$. We restrict attention to real-valued outcomes and treat $F$ as a random cumulative distribution function (cdf). We assume that $F$ is almost surely continuous, as this case is most relevant with regards to extremes and leads to a simpler presentation. However, generalizations are possible; see Remark 1.
@@ 2.2 Notions of tail calibration
To assess the calibration of forecasts for extreme events, we adopt the standard approach of extreme value theory and decompose $\operatorname{\mathsf{P}}(Y>x+t)=\operatorname{\mathsf{P}}(Y>t)\operatorname{\mathsf{P}}(Y>x+t\mid Y>t)$, for $x\geq 0$ and large $t$, i.e., $\operatorname{\mathsf{P}}(Y>t)$ positive but close to zero. Then, we separately consider calibration with respect to both factors on the right-hand side. Under auto-calibration,
@@
To avoid this scaling issue , we define the excess probability integral transform (PIT) as
@@ 3 Probabilistic tail calibration and peaks-over-threshold models
    The distribution of $Y$ is in the max-domain of attraction of a nondegenerate distribution.
Assumption (A1) is standard in extreme value theory, and by Balkema and de Haan (1974), it is equivalent to the existence of $\xi\in\mathbb{R}$ and a scaling function $\sigma_{Y}(\,\cdot\,)>0$ such that
@@
For the proof of Proposition 14, see the more general Proposition 31 in Appendix C in the supplement. If the scale parameter $\sigma_{F}(t)$ in (9) is allowed to be random, then $F$ can be probabilistically tail calibrated for $Y$ even when the GPD shape parameters $\xi$ and $\eta$ in (8) and (9) differ. For instance, if $\Delta\sim\Gamma(1/\gamma,1/\gamma)$ for some $\gamma>0$ and if, conditionally on $\Delta$, the distribution of $Y$ is $F=\operatorname{Exp}(\Delta)$, i.e., $\operatorname{\mathsf{P}}(Y\leq x\mid\Delta)=1-e^{-\Delta x}=F(x)$ for $x\geq 0$, then the unconditional distribution of $Y$ is $\operatorname{GPD}_{1,\gamma}$, so that (8) holds with $\xi=\gamma$, while $F_{t}$ is $\operatorname{Exp}(\Delta)$ too, and thus $F_{t}(x/\Delta)=1-e^{-x}$, which is (9) with random scale parameter $\sigma_{F}(t)=1/\Delta$ and with $\eta=0$.
Assumptions (A1), (A2), and property (a) of Proposition 14 are sufficient but not necessary conditions for probabilistic tail calibration. Indeed, according to Proposition 7, for non-random probabilistic forecasts, it suffices that $\operatorname{\mathsf{P}}(Y>t)/(1-F(t))\to 1$ as $t\to x_{Y}$.
@@ 4.1 Probabilistic tail calibration
Figure 5 displays the diagnostic plots for the tail calibration of the three forecasters using $n=10^{6}$ independent realizations of $\Delta$ and $Y$, see also Figure 10 in the supplement. The extremist forecast clearly has a heavier tail than the true distribution of the outcomes, and the nature of this miscalibration does not change as the threshold is increased.
@@ 4.2 Tail $\mathcal{B}$-calibration
To assess tail $\mathcal{B}$-calibration, the forecast-observation pairs $(F_{1},y_{1}),\dots,(F_{n},y_{n})$ can be stratified into $J\in\mathbb{N}$ mutually disjoint bins, $B_{1},\dots,B_{J}$, depending on variables representing the information contained in $\mathcal{B}$. The diagnostic plots for probabilistic tail calibration can then be applied separately to each bin of the forecasts and observations; if the forecasts appear probabilistically calibrated for all bins, then there is evidence to suggest that they are tail $\mathcal{B}$-calibrated. For example, the combined ratio in (10) can be calculated per bin,
@@
Displaying the diagnostic plots for each forecast method and each bin leads to a large number of plots to analyze. To simplify visualization, (tail) miscalibration can be summarized in a single value. This aligns with the general framework to assess conditional calibration proposed by Tsyplakov (2011) based on moment conditions and test functions, which is akin to how forecasts are compared in Giacomini and White (2006) and Nolde and Ziegel (2017). As a measure of tail calibration, consider the supremum distance $\sup_{u\in[0,1]}|\hat{R}_{t}(u)-u|$, which quantifies the maximum absolute difference between the combined ratio and the diagonal line in the diagnostic plot proposed above. Under probabilistic tail calibration, this difference should be close to zero for large $t$. To assess tail $\mathcal{B}$-calibration, $\hat{R}_{t}$ can be replaced by $\hat{R}_{t,j}$, and this distance can be calculated for $j=1,\dots,J$.
@@
for the ideal, climatological and extremist forecasters in Example 17 are shown as a function of the threshold in Figure 6, for $J=3$ bins for $\Delta$. For the ideal forecaster, the difference between $\hat{R}_{t,j}(u)$ and $u$ is essentially zero for all bins, confirming that it is tail $\sigma(\Delta)$-calibrated. The climatological forecaster, on the other hand, is clearly not calibrated conditionally on $\Delta$, despite being probabilistically tail and tail auto-calibrated. These plots for conditional calibration are shown for the empirical analogue (12) of the combined ratio in (4), though an analogous approach can serve to evaluate the occurrence ratio (5) or severity ratio (6).
@@ 5 Application
Three forecasting methods are compared. Firstly, we consider the 51-member ensemble forecasts generated by the ECMWF’s IFS ensemble. Ensemble forecasts are collections of point forecasts, forming an empirical predictive distribution function. An empirical distribution comprised of 51 sample values is unlikely to be useful when predicting extremes; in particular, the resulting forecast will assert that high thresholds will be exceeded with probability zero. We therefore also assess the performance of a smoothed version of the discrete ensemble forecast, which converts the IFS ensemble to a continuous forecast distribution by assuming that precipitation observations follow a censored logistic distribution, with location and scale parameters that are equal, respectively, to the mean and standard deviation of the corresponding IFS ensemble members. Finally, we present results for a statistical post-processing model that aims to remove systematic biases in the IFS ensemble to yield calibrated forecasts.
The post-processing method assumes that precipitation follows a certain parametric distribution, with location and scale parameters that are affine functions of the IFS ensemble mean and standard deviation, respectively. This simple approach is generally referred to as ensemble model output statistics in the post-processing literature (Gneiting et al., 2005). We implement this post-processing model with two choices of parametric distribution: a logistic distribution, which is commonly employed in studies of precipitation forecasts (see e.g. Messner et al., 2014); and a generalized extreme value (GEV) distribution, which has been proposed as a means to better forecast extreme outcomes (Lerch and Thorarinsdottir, 2013; Scheuerer, 2014). Both distributions are censored below at zero. The parameters of the post-processing models are estimated by minimizing the continuous ranked probability score (CRPS) over 20 years of twice-weekly reforecasts at the corresponding station and lead time; further details regarding the data can be found in Demaeyer et al. (2023).
[…]
@@
The poor tail calibration of the logistic post-processed forecasts is a deficiency of the chosen distribution. This is irrespective of the number of ensemble members: Even if we had no covariate information, we could still issue the climatological distribution of the outcome, which is auto-calibrated and therefore tail auto-calibrated. However, when employing existing forecast evaluation techniques, practitioners would generally conclude that this post-processing method generates reliable forecasts, which is misleading. In contrast, our diagnostic tools for tail calibration inform us that the forecast tails are too light, resulting in unreliable forecasts for extreme events. This information can be used to improve forecasts, for example by encouraging practitioners to experiment with alternative post-processing methods, based on different distributional assumptions.
[…]
@@ 6 Discussion
Tail calibration can be assessed with respect to different information sets, though we imagine the unconditional case will be of most interest in practical applications. To assess probabilistic tail calibration, we introduce graphical diagnostic tools that not only demonstrate whether or not a forecast is tail calibrated, but also elucidate what types of biases occur in the tail; for example, whether the predictive distributions are too heavy- or light-tailed. Formal testing procedures for tail calibration are an interesting avenue for future work, where the crux will be to choose a data-dependent threshold that is representative of the asymptotic regime. Such tests could either be classical tests for fixed pre-specified sample sizes, or monitoring procedures as introduced by Arnold et al. (2023), who propose a sequential procedure for testing probabilistic calibration based on $e$-processes, recognizing that forecasting is often an inherently sequential task.
@@
When the unconditional distribution of the outcome variable is in the max-domain of attraction of a non-degenerate distribution (Assumption (A1)), probabilistic tail calibration is equivalent to correctly forecasting the shape and scale parameters of the GPD. Checks for probabilistic tail calibration therefore serve as a more general framework with which to validate tail models, since they are also applicable to distributions not satisfying Assumption (A1). Future work may investigate whether notions of tail calibration can conversely be used to estimate quantities relevant for extreme value analysis, such as tail indices.
@@
Acknowledgments
We would like to thank two anonymous referees for their constructive feedback. Johan Segers gratefully acknowledges helpful questions and suggestions from participants, in particular Anja Janßen and Clément Dombry, of the EXSTA Kick-Off Workshop at Université Paris Cité, Laboratoire MAP5, June 5–7, 2024.
@@ Appendix A Omitted proofs and calculations
Write $t_{n}=x_{Y}-1/n$. We have $\mathbf{1}\{Y>t_{n}\}\to\mathbf{1}\{Y\geq x_{Y}\}=\mathbf{1}\{Y=x_{Y}\}$. By the conditional dominated convergence theorem (e.g. Kallenberg, 2002, Exercise 8 on page 117), $\operatorname{\mathsf{P}}(Y>t_{n}\mid\mathcal{B})\to\operatorname{\mathsf{P}}(Y=x_{Y}\mid\mathcal{B})$ as $n\to\infty$ a.s. Since $\operatorname{\mathsf{E}}[\operatorname{\mathsf{P}}(Y=x_{Y}\mid\mathcal{B})]=\operatorname{\mathsf{P}}(Y=x_{Y})>0$, we have $\operatorname{\mathsf{P}}(B)>0$ where $B=\{\operatorname{\mathsf{P}}(Y=x_{Y}\mid\mathcal{B})>0\}$.
Since $B\in\mathcal{B}$, we have
@@
$\frac{d}{du}\operatorname{\mathsf{P}}(Z_{F}\leq u)$ $=\frac{d}{du}\frac{1}{2}\{\operatorname{\mathsf{P}}(G_{+}(Y)\leq u)+\operatorname{\mathsf{P}}(G_{-}(Y)\leq u)\}$
@@
it follows that $z_{-}-1=z_{+}$, since $z\mapsto G(z)+G(z+1)$, $z\in[a-1,b]$ is strictly increasing, and $z_{+},z_{-}-1\in[a-1,b]$. Inserting this equality above yields $(d/du)\operatorname{\mathsf{P}}(Z_{F}\leq u)=1$, as required.